If a mass of 12 kg causes a spring to be extended from its equilibrium position by 1.2 mm, how far would the spring elongate if a 175 kg was used?
1
Expert's answer
2012-10-19T08:55:44-0400
According to the Hooke's law:

m*g = k*Δl ==> k = m*g/ΔL = 12[kg]*9.8[m/s²]/0.0012[m] = 98000 [kg/s²].

M*g = k*ΔL ==> ΔL = M*g/k = 175[kg]*9.8[m/s²]/98000[kg/s²] = 0.0175 m.

So, the elongation of the spring will be 1.75 cm.